Determination of the dissociation Constant (K_a) o
Answer #1

II.

pH of weakacid = 1/2(pka-logC)

c = concentration of weakacid = 0.1 M

pka = ?

   2.39 = 1/2(x-log0.1)

x = pka = 3.78

ka = 10^-pka

    ka = 10^-(3.78)

dissociation constant ( ka) = 1.66*10^-4

III.

A.

          HF(aq) <===> H^+(aq) + F^-(aq)

initial 0.5 M            0        0

at equil 0.5-x            x        x


ka = [H^+][F^-]/[HF]

7.1*10^-4 = x^2/(0.5-x)

x = [H^+]=[F^-] = 0.0185 M

[HF] = 0.5-0.018 = 0.482 M

pH = -log(H^+)

     = -log(0.0185)

    = 1.732

ka = cx^2

7.1*10^-4 = 0.5x^2

x = degree of dissociation = 0.0377

% of dissociation = 0.0377*100 = 3.77%